OpenStudy (anonymous):

How to change 1517 over 45 into a mixed number?

ganeshie8 (ganeshie8):

you familiar with long division ?

OpenStudy (anonymous):

1517 45 Divide the numerator, '1517' by the denominator, '45', to get 33 with remainder 32. The remainder is the new numerator. The quotient, '33', becomes the whole number in the mixed number. 33 32 45 The resulting mixed number is: 33 32 45

ganeshie8 (ganeshie8):

\(\frac{1517}{45} = \frac{33*45 + 32}{45}\) \( = \frac{33*45}{45} + \frac{32}{45}\) \( = 33 + \frac{32}{45}\) \( = 33 \frac{32}{45}\)

OpenStudy (anonymous):

^ ty :3

OpenStudy (anonymous):

Thanks

ganeshie8 (ganeshie8):

yw! :)